OpenStudy (anonymous):

If you can earn 5%, how much will you have to save each year if you want to retire in 35 years with $1 million? =(1.05)^(35) this is what i thought formula was but cannot seem to finish it. thanks in advance

OpenStudy (anonymous):

How much are you starting with? 5% of 0 = 0.

OpenStudy (anonymous):

thats exactly what it says in the book, i need to find the amount that i would have to start with

OpenStudy (anonymous):

Got it. Sorry, idiot moment.

OpenStudy (anonymous):

no worries my brain is fried also :)

OpenStudy (anonymous):

There is a formula for compounding interest, I just can't find it. Perhaps it's in your textbook or notes?

OpenStudy (anonymous):

i thought so but as yet i cannot find it

OpenStudy (anonymous):

11000.72 is what i came up with but cant remember how i got there :)

OpenStudy (anonymous):

A=75000(1+\{0.05}{1})^{1*15} i think i am missing something somewhere

OpenStudy (anonymous):

oops wrong formula for this question

OpenStudy (anonymous):

I don't see time anywhere...

OpenStudy (anonymous):

told you brain fried

OpenStudy (anonymous):

thanks anyway will have to keep messing around with formulas until i find the correct one

OpenStudy (anonymous):

just made adjustment to question

OpenStudy (anonymous):

\(A = P(1 + r)^t\) THis came from and about.com page - can't verify it's accuracy...

OpenStudy (anonymous):

=(1.05)^(35)

OpenStudy (anonymous):

I would say: 1 million = \(x(1.05)^{35}\) but I can't be sure.

OpenStudy (anonymous):

$1,000,000.00 35 5.516015368

OpenStudy (anonymous):

35 * 5.516015368 does not add up

OpenStudy (anonymous):

thats where i need help to finish this formula

OpenStudy (anonymous):

I'm looking at this page now... http://qrc.depaul.edu/studyguide2009/notes/savings%20accounts/compound%20interest.htm

OpenStudy (anonymous):

thx have just gone to link

OpenStudy (anonymous):

will give that a try when i find that calculator

OpenStudy (anonymous):

\(1.05^{35} = 5.51601536759226\) \(1,000,000 \div 5.51601536759226 = 181,290.285352577\) I don't know if this helps...